The rate of heat lost through a 3.5 inch thickness of insulation is 2200 Btu per hours. What is the rate of heat lost through a 5-inch thickness of the same insulation?
Answer by stanbon(75887)

The rate of heat lost through a 3.5 inch thickness of insulation is 2200 Btu per hours. What is the rate of heat lost through a 5-inch thickness of the same insulation?
---
thickness and loss of heat are inversely related.
---
h = k/t
---
Solve for "k" using "heat lost through a 3.5 inch thickness of insulation is 2200 Btu"
2200 = k/3.5
k = 3.5*2200 = 7700
----------------------------
Equation:
h = 7700/t
----
What is the rate of heat lost through a 5-inch thickness of the same insulation?
h = 7700/5 = 1540 Btu per hour
=================================
Cheers,
Stan H.